OpenStudy (anonymous):

Anyone taking the Algebra 2 A Final Exam? If so, i need help/ answers.

OpenStudy (anonymous):

All ready took it but I can help

OpenStudy (anonymous):

Is there anyway you can message me the answers? :/

OpenStudy (anonymous):

I'd need to know what your questions are first. Each semester exam has different questions

OpenStudy (anonymous):

Which equation illustrates the Associative Property of Addition? (2 points)

OpenStudy (anonymous):

For some reason it wont let me copy and paste the answers. but that is question #1

OpenStudy (anonymous):

here's an example of it: 5+(2+1)=(5+2)+1

OpenStudy (anonymous):

In other words, it's an addition equation where you can change the parenthesis around and it won't mess up the problem.

OpenStudy (anonymous):

Thanks, i got it! here is #8 Which of the following is the equation in slope-intercept form of the line through the points (1, 4) and (–1, –7)?

OpenStudy (anonymous):

I really do appreciate you helping me btw.

OpenStudy (anonymous):

well the slope is 11/2 so try that and you're welcome

OpenStudy (anonymous):

is it y=11/2x+3/2 or y=11/2x-3/2 ?

OpenStudy (anonymous):

y=11/2x-3/2

OpenStudy (anonymous):

thank God for graphing utilities it's so much faster

OpenStudy (anonymous):

Got it! here is #9 Which of the following is the equation in slope-intercept form of the line through the points (3, 1) and (–5, 2)?

OpenStudy (anonymous):

slope is -1/8 should be something like y=-1/8x+4/3

OpenStudy (anonymous):

not sure about the actual second fraction though but it would be a plus sign

OpenStudy (anonymous):

Which statement is true about the line with equation y=-1/3x+2 ? (2 points) The x-intercept is (–6, 0). The y-intercept is (0, –2). The standard equation is x+3y=6 . The point (1,2/3) is on the line.

OpenStudy (anonymous):

The standard equation is x+3y=6

OpenStudy (anonymous):

How can you translate the graph of y=x to obtain the graph of y= (x-4)+2? (2 points) 4 units left and 2 units up 4 units left and 2 units down 4 units right and 2 units up 4 units right and 2 units down

OpenStudy (anonymous):

4 units right and 2 up

OpenStudy (anonymous):

Paul’s Plumbing charges $55 to visit a job site plus $35 per hour for labor. Reliable Plumbing charges $75 to visit a job site plus $30 per hour for labor. For how many hours of labor would the total charge be the same for the two companies? (2 points) 2 hours 4 hours 15 hours 26 hours

OpenStudy (anonymous):

4 hours

OpenStudy (anonymous):

Which point lies in the xz-plane and below the xy-plane? (2 points) (2, 0, –4) (0, 2, –4) (–2, 0, 4) (0, –2, 4)

OpenStudy (anonymous):

Im skipping alot of questions cause there are graphs ans what not. :c

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

i think its a?

OpenStudy (anonymous):

actually I don't know that one

OpenStudy (anonymous):

Which of the following is a quadratic function in standard form for the set of points(0,-1) (,3 (2,9) ?  

OpenStudy (anonymous):

Sorry i meant, Which of the following is a quadratic function in standard form for the set of points(0,-1) (1,3) (2,9) ?

OpenStudy (anonymous):

what are your answers for this one?

OpenStudy (anonymous):

F(x)=x^2+3x-1 F(x)=x^2+5x-1 F(x)=-x^2+5x-1 F(x)=3x^2+x-1

OpenStudy (anonymous):

F(x)=x^2+3x-1

OpenStudy (anonymous):

sorry if I'm a little slow at answering. I'm doing English at the same time. LOL

OpenStudy (anonymous):

Its all good man!

OpenStudy (anonymous):

Which is the correct description of the graph of the function y=-2x^2+x?

OpenStudy (anonymous):

what are your answers? The graph is a parabola that opens downward. So probably something like that

OpenStudy (anonymous):

Is anyone doing English 3 highschool from ashworth?

OpenStudy (anonymous):

I have two answers that say downward, which are there. #1 (1/4),(1,8) #2(-1,4), (-2/8)

OpenStudy (anonymous):

@tobie_quench I'm doing English 3 from FLVS but I might be able to help

OpenStudy (anonymous):

what's the name of your book?

OpenStudy (anonymous):

English Book?

OpenStudy (anonymous):

#2 I think

OpenStudy (anonymous):

@tobie_quench To Kill a Mocking Bird and The House On Mango Street

OpenStudy (anonymous):

@tobie_quench oh wait you didn't mean for reading did you?

OpenStudy (anonymous):

@tobie_quench uh for non-reading all of my stuff is online no real books

OpenStudy (anonymous):

Me too... that is what makes everything harder!

OpenStudy (anonymous):

@tobie_quench really? I find it so much easier.

OpenStudy (anonymous):

@tobie_quench but that's because I can go at my own pace. I'm WAY ahead now. 15 and in 11th grade.

OpenStudy (anonymous):

congrats to you! im a good bit slower

OpenStudy (anonymous):

Hey man lol

OpenStudy (anonymous):

I say it all depends on the person. online schooling is better because it's more suitable for the slower and faster people

OpenStudy (anonymous):

Yay!!! 3-person conversation

OpenStudy (anonymous):

I just have a hard time staying focused

OpenStudy (anonymous):

I'm random. -_-'

OpenStudy (anonymous):

I have ADHD. I know what it's like

OpenStudy (anonymous):

I smoke alot, i know whats its like as well.

OpenStudy (anonymous):

and the readings, I just don't understand or like reading them!

OpenStudy (anonymous):

But I also have Asperger's Syndrome (means I'm smart and different sort of) and OCD (Obsessive Compulsive Disorder)

OpenStudy (anonymous):

I promised my self that I would never smoke or do any drugs. I'm not even going to get a tattoo. (Unless I decide to get a pair of Dragon wings on my back)

OpenStudy (anonymous):

I always told myself the same thing, now i love smoking. But im trying my best to stay sober, i havent smoked in 3 weeks my friend!

OpenStudy (anonymous):

That's a great start!!

OpenStudy (anonymous):

I just got my first fan! YAY! thanks

OpenStudy (anonymous):

You're welcome!!! =)

OpenStudy (anonymous):

Haha thanks man.

OpenStudy (anonymous):

I like fanning new friends!!!

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

Haha me too!

OpenStudy (anonymous):

although I fanned @nubeer just because their profile pic is Itachi from Naruto/Naruto Shippuden

OpenStudy (anonymous):

Itachi's name is awesome!!! It means weasel!! LOL

OpenStudy (anonymous):

hey i have 3 questions left, can you help me finish? :]

OpenStudy (anonymous):

of course!!!

OpenStudy (anonymous):

Thanks dude, i really appreciate it!

OpenStudy (anonymous):

Sorry i meant, Which of the following is a quadratic function in standard form for the set of points(0,-1) (1,3) (2,9) ?

OpenStudy (anonymous):

F(x)=x^2+3x-1 F(x)=x^2+5x-1 F(x)=-x^2+5x-1 F(x)=3x^2+x-1

OpenStudy (anonymous):

Dude sorry that wasnt the right one! My fault

OpenStudy (anonymous):

Which is the correct description of the graph of the function y=-2x^2+x?

OpenStudy (anonymous):

I have two answers that say downward, which are there. #1 (1/4),(1,8) #2(-1,4), (-2/8)

OpenStudy (anonymous):

#2

OpenStudy (anonymous):

Write the polynomial function f(t)=-t-t(2t)^2 in standard form. Classify the function by degree and number of terms

OpenStudy (anonymous):

Not sure about this one. Do you have answers?

OpenStudy (anonymous):

or is it an essay question?

OpenStudy (anonymous):

Yes i do, give me a sec lol

OpenStudy (anonymous):

i have to type them since its not letting me cope and paste

OpenStudy (anonymous):

that's fine

OpenStudy (anonymous):

f(t)=4t^2-2t quadratic binomial f(t)=4t^2-2t quadratic trinomial f(t)=-4t^3-t cubic binomial f(t)=-4t^3-t cubic trinomial

OpenStudy (anonymous):

f(t)=-4t^3-t cubic binomial

OpenStudy (anonymous):

Okay I have time for one more but then I have to go okay

OpenStudy (anonymous):

last one man. Which of the following are solutions to the equation x^2+2x+2=0 ?

OpenStudy (anonymous):

answers?

OpenStudy (anonymous):

-2,-1 -1sq root of 3,-1, -sqrt 3 -1+i,-1+i 1+i,1-i

OpenStudy (anonymous):

Sorry for the trouble bro!

OpenStudy (anonymous):

hey, im taking algebra 2 right now and was wondering if any of you could please help